# Thermodynamically consistent Cahn-Hilliard-Navier-Stokes equations using the metriplectic dynamics formalism

### v2

@@ Abstract
Cahn-Hilliard-Navier-Stokes (CHNS) systems describes flows with two-phases, e.g., a liquid with bubbles. Obtaining constitutive relations for general dissipative processes for such a systems, which are thermodynamically consistent, can be a challenge. We show how the metriplectic 4-bracket formalism (Morrison and Updike, 2023) achieves this in a straightforward, in fact algorithmic, manner. First, from the noncanonical Hamiltonian formulation for the ideal part of a CHNS system we obtain an appropriate Casimir to serve as the entropy in the metriplectic formalism that describes the dissipation (e.g. viscosity, heat conductivity and diffusion effects). General thermodynamics with the thermodynamic conjugates of concentration and chemical potential are included. Having expressions for the Hamiltonian (energy), entropy, and Poisson bracket, we describe a procedure for obtaining a metriplectic 4-bracket that describes thermodynamically consistent dissipative effects. The 4-bracket formalism leads naturally to a general CHNS system that allows for anisotropic surface energy effects. This general CHNS system reduces to cases in the literature, to which we can compare.
@@ 1 Introduction
The well-known Navier-Stokes equations govern the motion of a single-phase fluid. However, in the case of two-phase fluids, chemical reactions, changes of phase, and migration between substances of phases become significant and cannot be disregarded. J. W.  Cahn and J. E. Hilliard were the first to formulate the mathematical equations that describe phase separation in a such a binary fluid (Cahn and Hilliard, 1958). Here we investigate generalizations that combine the Cahn-Hilliard equation with equations that describe the dynamics of fluid flow, referred to as Cahn-Hilliard-Navier-Stokes (CHNS) systems. CHNS systems aim to describe the hydrodynamic properties of a mixture of two phases such as bubbles in a liquid. To narrow down the already broad scope, we assume that the two fluids share the same velocity field, yet we allow for both extended thermodynamics and diffusive interfaces between the two phases.
@@
The algorithm has four steps: i) Select a set of dynamical variables. For a CHNS system these will be $\psi\coloneqq\{\mathbf{m}=\rho\mathbf{v},\rho,\tilde{c}=\rho c,\sigma=\rho s\}$, which are the momentum density, mass density, volume concentration of one of the constituents, and entropy density, respectively. ii) The next step is to select energy and entropy functionals, $H$ and $S$, dependent on the dynamical variables. The choice of these functionals is based on the physics of the phenomena one wishes to describe. iii) The third step of the algorithm is to obtain the noncanonical Poisson bracket (see Morrison, 1998) of the ideal (nondissipative) part of the theory that has the chosen entropy as a Casimir invariant. Since the work of Morrison and Greene (1980), Poisson brackets for a great many systems, including fluid and magnetofluid systems, have been found (e.g. Morrison, 1982; Thiffeault and Morrison, 2000; Abdelhamid et al., 2015; D’Avignon et al., 2016; Zaidni et al., 2023). Thus, this step may be immediate. Alternatively, it may be achieved by a coordinate change from a known Hamiltonian theory in order to align with the chosen entropy functional. In either case, we obtain at this stage a noncanonical Hamiltonian system. iv) The final step is to construct a metriplectic 4-bracket as described in Morrison and Updike (2023). Although there are standard metriplectic 4-bracket constructions, there is freedom at this last step to describe a variety of types of dissipation. However, a natural choice follows upon consideration of the form of an early metriplectic bracket (Morrison, 1984b). Given $H$, $S$, and the 4-bracket, the dynamical system with thermodynamically consistent dissipation is produced.
@@
We apply the algorithm to two cases. First, in § 2, we consider a system where the fluid thermodynamics is extended by allowing the internal energy to depend on a concentration variable, with the chemical potential being its thermodynamic dual. Because Gibbs introduced the notion of chemical potential, we refer to the Hamiltonian version of this fluid systems as the Gibbs-Euler (GE) system and the dissipative version as the Gibbs-Navier-Stokes (GNS) system. It is a thermodynamically consistent version of the compressible Navier-Stokes equations with the inclusion of this concentration variable for describing a second phase of the fluid. The GNS system generalizes the early work of Eckart (1940a, b) and the treatment in (de Groot and Mazur, 1962); it allows for all possible thermodynamics fluxes. Next, in § 3, a general form of CHNS system is produced, a form that models surface tension effects and allows for diffuse interfaces. Our work is motivated in large part by the substantial works of Anderson et al. (2000) and Guo and Lin (2015), which we generalize by obtaining a class of systems that includes theirs as special cases. There is a huge literature on this topic and these papers contain many important references to previous work. (Also, see ten Eikelder et al., 2023, for a recent review.)
@@
The GNS system of § 2 serves as a straightforward example of our algorithm. In §§ 2.1 we describe the set of dynamical variables, properties of the system, the energy and entropy functionals $H$ and $S$. This amounts to the first and second steps of the algorithm. Then in §§ 2.2 the Hamiltonian formulation of the dissipation free part of the system is presented. This is the third step of the algorithm where the Poisson bracket is obtained, after a brief review of the noncanonical Hamiltonian formalism. Given the early work of Morrison and Greene (1980) and the classification of extensions in (Thiffeault and Morrison, 2000), this step is immediate. Based on the early and recent works of Morrison (1984b) and Morrison and Updike (2023) the fourth step of the algorithm is also immediate. In §§ 2.3 we first review the metriplectic 4-bracket formalism and present the realization that applies for the GNS system. Thus, the thermodynamically consistent GNS system is determined. In §§ 2.4 we obtain the metriplectic 2-bracket equations of motion, and the determined fluxes and affinities, making connection to standard irreversible thermodynamics. Using the results of § 2, we proceed in § 3 to obtain the main result of the paper, our general CHNS system that can describe diffuse interface effects. The first and second steps of our algorithm are taken in §§ 3.1, while the third step, obtaining the correct Poisson bracket, is undertaken in §§ 3.2. In order to complete this step, one must find the Poisson bracket for which the entropy of the second step is a Casimir invariant, which we find can be achieved by a simple coordinate transformation. The fourth step of the algorithm is taken in §§ 3.3. Here a choice of metriplectic 4-bracket gives a general class of thermodynamically consistent CHNS systems, a class that contains previous results as special cases. The formalism also shows how one can transform to a simple entropy variable at the expense of a more complicated internal energy. As in § 2, in §§ 3.4 we reduce to the metriplectic 2-bracket. Finally in § 4 we briefly summarize and make a few comments about ongoing and future work.
@@ 2.1 Description of the Gibbs-Navier-Stokes system
The GNS for 2 phase flow proceeds on familiar ground (Eckart, 1940a, b; de Groot and Mazur, 1962). It amounts to the single phase thermodynamic Navier-Stokes system or as it is sometimes called the Fourier Navier-Stokes system with the dispersed phase described by the addition of a concentration variable, $c$, giving the set of dynamical variables $\psi=\{\mathbf{v},\rho,c,s\}$. Here we review global aspects of this known system, before showing how it emerges from the metriplectic formalism.
@@ 2.2 Noncanonical Poisson bracket of the Gibbs-Euler system
Given that the mixture is assumed to be confined in the domain $\Omega$, the Eulerian scalars (volume forms) $(\rho,\tilde{c},\sigma)$ are functions from space-time $\Omega\mapsto\mathbb{R}\rightarrow\mathbb{R}$, while the vector field $\mathbf{m}$ maps $\Omega\times\mathbb{R}\mapsto T\Omega$, where $T\Omega$ stands for the tangent bundle of the manifold $\Omega$. We will forgo formal geometric considerations and suppose our infinite-dimensional phase space has coordinates $\psi=(\mathbf{m},\rho,\tilde{c},\sigma)$ and observables are functionals that map $\psi\mapsto\mathbb{R}$ at each fixed time. We will denote the space of such functionals by $\mathcal{B}$. Then a Poisson bracket is a bilinear operator $\mathcal{B}\times\mathcal{B}\mapsto\mathbb{R}$ that fulfills the Leibniz rule and is a realization of a Lie algebra (see e.g. Sudarshan and Mukunda, 1974, chap. 14). The Leibniz rule follows from that for the variational or functional derivative of $F\in\mathcal{B}$, defined by
$\delta F[\psi;\delta\psi)=\lim_{\epsilon\rightarrow 0}\frac{F(\psi+\epsilon\delta\psi)-F(\psi)}{\epsilon}=\left\langle\frac{\delta F}{\delta\psi},\delta\psi\right\rangle\,,$
[…]
@@
where we compactified our notation by defining $F_{\mathbf{m}}\!\!\coloneqq\delta F/\delta\mathbf{m}$, $F_{\rho}\!\!\coloneqq\delta F/\delta\rho$, etc., the functional derivatives with respect to the various coordinates $\psi$. That this is the appropriate Poisson bracket is immediate; it is the Lie-Poisson bracket originally given by Morrison and Greene (1980) with the addition of the last line of (15) involving the concentration, another volume density variable $\tilde{c}$. Adding such a dynamical variable is common place in the fluid modeling of plasmas over the last decades and fits within the general theory for extension given by Thiffeault and Morrison (2000). By construction we have a Poisson bracket that is a bilinear, antisymmetric, and either by the extension theory or a relatively easy direct calculation using the techniques of Morrison (1982) it can be shown to satisfy the Jacobi identity, i.e.,
@@
lies in our set of Casimirs. This quantity was first shown to be a Casimir for the ideal fluid in (Morrison, 1982) and used for the thermodynamically consistent Navier-Stokes metriplectic system in (Morrison, 1984b). We note in passing, for other theories that might have a nontraditional dynamical equilibrium playing the role of thermodynamic equilibrium, one may wish to choose another Casimir.
@@ 2.3 Metriplectic 4-bracket for the Gibbs-Navier-Stokes system
Now let us turn to our fourth and final step of the algorithm, construction of the metriplectic 4-bracket. To this end we review the formalism of Morrison and Updike (2023) in general terms and then apply it to the GNS system, which is a generalization of an example given in that work.
@@ 2.3.1 General metriplectic 4-bracket dynamics
The metriplectic 4-bracket theory was introduced by Morrison and Updike (2023) to describe the dissipative dynamic. Let us briefly recall the metriplectic 4-bracket description in infinite dimensions. In this description, we consider the dynamics of classical field theories with multi-component fields
@@
The minimal metriplectic properties of metriplectic 4-brackets are reminiscent of the algebraic properties possessed by a curvature tensor. In fact, every Riemannian manifold naturally has a metriplectic 4-bracket, and $(S,H;S,H)$ provides a notion of sectional curvature (see Morrison and Updike, 2023).
@@
For the full metriplectic dynamics we would add the Poisson bracket contribution to the above. Equation (42) is written so as to show that it amounts to a gradient systems with the entropy $S$ as generator.
@@ 2.3.2 General Kulkarni-Nomizu construction
We can easily create specific metriplectic 4-brackets that have the minimal metriplectic properties: the requisite symmetries and the positive semidefiniteness $(S,H;S,H)$. We do this by using the Kulkarni-Nomizu (K-N) product (Kulkarni, 1972; Nomizu, 1971). See also Fiedler (2003) for relevant theorems. Consistent with the bracket formulation of (33), we deviate from the conventional K-N product by working on the dual. Given two symmetric operator fields, say $\sum$ and $M$, operating on the variational derivatives; we again use the subscript notation when convenient,
@@
which could follow from (46) if we added an additional argument to $\Sigma$. Then, with a corresponding form for $M$ the algebraic curvature symmetries would be induced in the integrand. This is the case for our present purposes, where we assume the specific K-N form given in Morrison and Updike (2023), viz.  where the 4-bracket is given by
@@ 2.4 GNS metriplectic 2-bracket and conventional fluxes and affinities
For completeness we demonstrate two things in this subsection: how the metriplectic 4-bracket formalism relates to the original binary metriplectic formalism given in Morrison (1984a, b, 1986) and how it relates to conventional nonequilibrium thermodynamics, making the connection between the 4-bracket K-N construction and the phenomenology of thermodynamics fluxes and affinities (sometimes called thermodynamic forces).
@@ 2.4.1 GNS metriplectic 2-bracket
As noted above we are concerned with the metriplectic dynamics introduced in Morrison (1984a, b, 1986) (see also Morrison, 2009; Coquinot and Morrison, 2020), but we mention that other binary brackets for describing dissipation were presented over the years (e.g. Kaufman and Morrison, 1982; Kaufman, 1984; Morrison and Hazeltine, 1984; Grmela, 1984; Grmela and Öttinger, 1997; Beris and Edwards, 1994; Edwards, 1998). In addition we mention a recent alternative approach to multiphase fluids, one based on constrained variational principles, that is given in Eldred and Gay-Balmaz (2020). We refer the reader to Morrison and Updike (2023) for comparisons with other formulations and how they emerge from the metriplectic 4-bracket.
@@
which follows because $\{S,\mathcal{F}\}\equiv 0$ and $({S},{H})_{H}\equiv 0$. As shown in Morrison and Updike (2023) the metriplectic 2-bracket emerges from the 4-bracket as follows:
@@
where the total energy density is used instead of the entropy density as one of our dynamical variables. That this is possible is well known in thermodynamics because the entropy must be a monotonic increasing function of the internal energy, which allows via the inverse function theorem transformation between the extensive energy or extensive entropy representations (Callen, 1966). We will denote these density variables in order by $\xi_{\alpha}$, $\alpha=1,\dots,6$. With this choice the Hamiltonian (18) is given by
@@
where without loss of generality we set $A^{66}=1$ and $B^{\alpha\beta}=L_{\alpha\beta}$ for this case. In §§ 2.4.2 we will follow Coquinot and Morrison (2020) and show how the bracket of (76) fits into the framework of conventional nonequilibrium thermodynamics as, e.g., described in de Groot and Mazur (1962). In this way we will physically identify the meaning of $L_{\alpha\beta}$.

### v3

@@ Abstract
Cahn-Hilliard-Navier-Stokes (CHNS) systems describes flows with two-phases, e.g., a liquid with bubbles. Obtaining constitutive relations for general dissipative processes for such a systems, which are thermodynamically consistent, can be a challenge. We show how the metriplectic 4-bracket formalism [30] achieves this in a straightforward, in fact algorithmic, manner. First, from the noncanonical Hamiltonian formulation for the ideal part of a CHNS system we obtain an appropriate Casimir to serve as the entropy in the metriplectic formalism that describes the dissipation (e.g. viscosity, heat conductivity and diffusion effects). General thermodynamics with the concentration variable and its thermodynamics conjugate, the chemical potential, are included. Having expressions for the Hamiltonian (energy), entropy, and Poisson bracket, we describe a procedure for obtaining a metriplectic 4-bracket that describes thermodynamically consistent dissipative effects. The 4-bracket formalism leads naturally to a general CHNS system that allows for anisotropic surface energy effects. This general CHNS system reduces to cases in the literature, to which we can compare.
@@ 1 Introduction
The well-known Navier-Stokes equations govern the motion of a single-phase fluid. However, in the case of two-phase fluids, chemical reactions, changes of phase, and migration between substances of phases become significant and cannot be disregarded. J. W.  Cahn and J. E. Hilliard were the first to formulate the mathematical equations that describe phase separation in a such a binary fluid [4]. Here we investigate generalizations that combine the Cahn-Hilliard equation with equations that describe the dynamics of fluid flow, referred to as Cahn-Hilliard-Navier-Stokes (CHNS) systems. CHNS systems aim to describe the hydrodynamic properties of a mixture of two phases such as bubbles in a liquid. To narrow down the already broad scope, we assume that the two fluids share the same velocity field, yet we allow for both extended thermodynamics and diffusive interfaces between the two phases.
@@
The algorithm has four steps: i) Select a set of dynamical variables. For a CHNS system these will be $\psi\coloneqq\{\mathbf{m}=\rho\mathbf{v},\rho,\tilde{c}=\rho c,\sigma=\rho s\}$, which are the momentum density, mass density, volume concentration of one of the constituents, and entropy density, respectively. ii) The next step is to select energy and entropy functionals, $H$ and $S$, dependent on the dynamical variables. The choice of these functionals is based on the physics of the phenomena one wishes to describe. iii) The third step of the algorithm is to obtain the noncanonical Poisson bracket [see 25] of the ideal (nondissipative) part of the theory that has the chosen entropy as a Casimir invariant. Since the work of Morrison and Greene [28], Poisson brackets for a great many systems, including fluid and magnetofluid systems, have been found [e.g. 21, 36, 1, 7, 38]. Thus, this step may be immediate. Alternatively, it may be achieved by a coordinate change from a known Hamiltonian theory in order to align with the chosen entropy functional. In either case, we obtain at this stage a noncanonical Hamiltonian system. iv) The final step is to construct a metriplectic 4-bracket as described in Morrison and Updike [30]. Although there are standard metriplectic 4-bracket constructions, there is freedom at this last step to describe a variety of types of dissipation. However, a natural choice follows upon consideration of the form of an early metriplectic bracket [23]. Given $H$, $S$, and the 4-bracket, the dynamical system with thermodynamically consistent dissipation is produced.
@@
We apply the algorithm to two cases. First, in § 2, we consider a system where the fluid thermodynamics is extended by allowing the internal energy to depend on a concentration variable, with the chemical potential being its thermodynamic dual. Because Gibbs introduced the notion of chemical potential, we refer to the Hamiltonian version of this fluid systems as the Gibbs-Euler (GE) system and the dissipative version as the Gibbs-Navier-Stokes (GNS) system. It is a thermodynamically consistent version of the compressible Navier-Stokes equations with the inclusion of this concentration variable for describing a second phase of the fluid. The GNS system generalizes the early work of Eckart [8, 9] and the treatment in [16]; it allows for all possible thermodynamics fluxes. Next, in § 3, a general form of CHNS system is produced, a form that models surface tension effects and allows for diffuse interfaces. Our work is motivated in large part by the substantial works of Anderson et al. [2] and Guo and Lin [17], which we generalize by obtaining a class of systems that includes theirs as special cases. There is a huge literature on this topic and these papers contain many important references to previous work. [Also, see 11, for a recent review.]
@@
The GNS system of § 2 serves as a straightforward example of our algorithm. In §§ 2.1 we describe the set of dynamical variables, properties of the system, the energy and entropy functionals $H$ and $S$. This amounts to the first and second steps of the algorithm. Then in §§ 2.2 the Hamiltonian formulation of the dissipation free part of the system is presented. This is the third step of the algorithm where the Poisson bracket is obtained, after a brief review of the noncanonical Hamiltonian formalism. Given the early work of Morrison and Greene [28] and the classification of extensions in [36], this step is immediate. Based on the early and recent works of Morrison [23] and [30] the fourth step of the algorithm is also immediate. In §§ 2.3 we first review the metriplectic 4-bracket formalism and present the realization that applies for the GNS system. Thus, the thermodynamically consistent GNS system is determined. In §§ 2.4 we obtain the metriplectic 2-bracket equations of motion, and the determined fluxes and affinities, making connection to standard irreversible thermodynamics. Using the results of § 2, we proceed in § 3 to obtain the main result of the paper, our general CHNS system that can describe diffuse interface effects. The first and second steps of our algorithm are taken in §§ 3.1, while the third step, obtaining the correct Poisson bracket, is undertaken in §§ 3.2. In order to complete this step, one must find the Poisson bracket for which the entropy of the second step is a Casimir invariant, which we find can be achieved by a simple coordinate transformation. The fourth step of the algorithm is taken in §§ 3.3. Here a choice of metriplectic 4-bracket gives a general class of thermodynamically consistent CHNS systems, a class that contains previous results as special cases. The formalism also shows how one can transform to a simple entropy variable at the expense of a more complicated internal energy. As in § 2, in §§ 3.4 we reduce to the metriplectic 2-bracket. Finally in § 4 we briefly summarize and make a few comments about ongoing and future work.
@@ 2.1 Description of the Gibbs-Navier-Stokes system
The GNS for 2 phase flow proceeds on familiar ground [8, 9, 16]. It amounts to the single phase thermodynamic Navier-Stokes system or as it is sometimes called the Fourier Navier-Stokes system with the dispersed phase described by the addition of a concentration variable, $c$, giving the set of dynamical variables $\psi=\{\mathbf{v},\rho,c,s\}$. Here we review global aspects of this known system, before showing how it emerges from the metriplectic formalism.
@@ 2.2 Noncanonical Poisson bracket of the Gibbs-Euler system
Given that the mixture is assumed to be confined in the domain $\Omega$, the Eulerian scalars (volume forms) $(\rho,\tilde{c},\sigma)$ are functions from space-time $\Omega\mapsto\mathbb{R}\rightarrow\mathbb{R}$, while the vector field $\mathbf{m}$ maps $\Omega\times\mathbb{R}\mapsto T\Omega$, where $T\Omega$ stands for the tangent bundle of the manifold $\Omega$. We will forgo formal geometric considerations and suppose our infinite-dimensional phase space has coordinates $\psi=(\mathbf{m},\rho,\tilde{c},\sigma)$ and observables are functionals that map $\psi\mapsto\mathbb{R}$ at each fixed time. We will denote the space of such functionals by $\mathcal{B}$. Then a Poisson bracket is a bilinear operator $\mathcal{B}\times\mathcal{B}\mapsto\mathbb{R}$ that fulfills the Leibniz rule and is a realization of a Lie algebra [see e.g. 33, chap. 14]. The Leibniz rule follows from that for the variational or functional derivative of $F\in\mathcal{B}$, defined by
$\delta F[\psi;\delta\psi)=\lim_{\epsilon\rightarrow 0}\frac{F(\psi+\epsilon\delta\psi)-F(\psi)}{\epsilon}=\int_{\Omega}\frac{\delta F}{\delta\psi}\delta\psi,$
where $\delta F/\delta\psi$ is the functional derivative. This expression can be viewed as the directional derivative of a functional $F$ at $\psi$ in the direction $\delta\psi$ [see, e.g., 25, for a formal review of these notions].
@@
where we compactified our notation by defining $F_{\mathbf{m}}\!\!\coloneqq\delta F/\delta\mathbf{m}$, $F_{\rho}\!\!\coloneqq\delta F/\delta\rho$, etc., the functional derivatives with respect to the various coordinates $\psi$. That this is the appropriate Poisson bracket is immediate; it is the Lie-Poisson bracket originally given by Morrison and Greene [28] with the addition of the last line of (15) involving the concentration, another volume density variable $\tilde{c}$. Adding such a dynamical variable is common place in the fluid modeling of plasmas over the last decades and fits within the general theory for extension given by Thiffeault and Morrison [36]. By construction we have a Poisson bracket that is a bilinear, antisymmetric, and either by the extension theory or a relatively easy direct calculation using the techniques of Morrison [21] it can be shown to satisfy the Jacobi identity, i.e.,
@@
lies in our set of Casimirs. This quantity was first shown to be a Casimir for the ideal fluid in [21] and used for the thermodynamically consistent Navier-Stokes metriplectic system in [23]. We note in passing, for other theories that might have a nontraditional dynamical equilibrium playing the role of thermodynamic equilibrium, one may wish to choose another Casimir.
@@ 2.3 Metriplectic 4-bracket for the Gibbs-Navier-Stokes system
Now let us turn to our fourth and final step of the algorithm, construction of the metriplectic 4-bracket. To this end we review the formalism of [30] in general terms and then apply it to the GNS system, which is a generalization of an example given in that work.
@@ 2.3.1 General metriplectic 4-bracket dynamics
The metriplectic 4-bracket theory was introduced by Morrison and Updike [30] to describe the dissipative dynamic. Let us briefly recall the metriplectic 4-bracket description in infinite dimensions. In this description, we consider the dynamics of classical field theories with multi-component fields
@@
The minimal metriplectic properties of metriplectic 4-brackets are reminiscent of the algebraic properties possessed by a curvature tensor. In fact, every Riemannian manifold naturally has a metriplectic 4-bracket, and $(S,H;S,H)$ provides a notion of sectional curvature [see 30].
@@
For the full metriplectic dynamics we would add the Poisson bracket contribution to the above. Equation (42) is written so as to show that it amounts to a gradient system with the entropy $S$ as generator.
@@ 2.3.2 General Kulkarni-Nomizu construction
We can easily create specific metriplectic 4-brackets that have the minimal metriplectic properties: the requisite symmetries and the positive semidefiniteness $(S,H;S,H)$. We do this by using the Kulkarni-Nomizu (K-N) product [20, 31]. See also Fiedler [13] for relevant theorems. Consistent with the bracket formulation of (33), we deviate from the conventional K-N product by working on the dual. Given two symmetric operator fields, say $\sum$ and $M$, operating on the variational derivatives; we again use the subscript notation when convenient,
@@
which could follow from (46) if we added an additional argument to $\Sigma$. Then, with a corresponding form for $M$ the algebraic curvature symmetries would be induced in the integrand. This is the case for our present purposes, where we assume the specific K-N form given in [30], viz.  where the 4-bracket is given by
@@ 2.4 GNS metriplectic 2-bracket and conventional fluxes and affinities
For completeness we demonstrate two things in this subsection: how the metriplectic 4-bracket formalism relates to the original binary metriplectic formalism given in Morrison [22, 23, 24] and how it relates to conventional nonequilibrium thermodynamics, making the connection between the 4-bracket K-N construction and the phenomenology of thermodynamics fluxes and affinities (sometimes called thermodynamic forces).
@@ 2.4.1 GNS metriplectic 2-bracket
As noted above we are concerned with the metriplectic dynamics introduced in Morrison [22, 23, 24] [see also 26, 6], but we mention that other binary brackets for describing dissipation were presented over the years [e.g. 19, 18, 29, 14, 15, 3, 10]. In addition we mention a recent alternative approach to multiphase fluids, one based on constrained variational principles, that is given in Eldred and Gay-Balmaz [12]. We refer the reader to Morrison and Updike [30] for comparisons with other formulations and how they emerge from the metriplectic 4-bracket.
@@
which follows because $\{S,\mathcal{F}\}\equiv 0$ and $({S},{H})_{H}\equiv 0$. As shown in Morrison and Updike [30] the metriplectic 2-bracket emerges from the 4-bracket as follows:
@@
where the total energy density is used instead of the entropy density as one of our dynamical variables. That this is possible is well known in thermodynamics because the entropy must be a monotonic increasing function of the internal energy, which allows via the inverse function theorem transformation between the extensive energy or extensive entropy representations [5]. We will denote these density variables in order by $\xi_{\alpha}$, $\alpha=1,\dots,6$. With this choice the Hamiltonian (18) is given by
@@
where without loss of generality we set $A^{66}=1$ and $B^{\alpha\beta}=L_{\alpha\beta}$ for this case. In §§ 2.4.2 we will follow Coquinot and Morrison [6] and show how the bracket of (76) fits into the framework of conventional nonequilibrium thermodynamics as, e.g., described in de Groot and Mazur [16]. In this way we will physically identify the meaning of $L_{\alpha\beta}$.